OpenStudy (anonymous):

Complete the square to solve the quadratic equation x^2-7x+12=0 I have the answer but its saying its 3,4 i got x=7/2±√1/4

OpenStudy (ash2326):

@yomamabf We have here \[x^2-7x+12=0\] OK tell me How did you do this problem?

OpenStudy (anonymous):

This is a complete the square method so I did 7/2 times 7/2 i followed all the directions i got it but the only problem is why it came up as 3,4

OpenStudy (ash2326):

You have done right, Let's try that again \[x^2-7x+12+\frac{49}{4}-\frac{49}{4}=0\] What did you next?

OpenStudy (anonymous):

12 goes on the other side as -12 and added to 49/4 which comes out as 1/4 then it looks like x=7/2±√1/4

OpenStudy (anonymous):

why is it 3,4

OpenStudy (ash2326):

Oh you have the answer in front of you:) What's the value of this \[\sqrt{\frac{1}{4}}=\frac{1}{\sqrt 4}=????\]

OpenStudy (anonymous):

? what ......

OpenStudy (anonymous):

are you sure you know what you're doing

OpenStudy (anonymous):

nvm dude omfg i got it i watched a youtube video omfg

OpenStudy (ash2326):

Yeah \[\sqrt{\frac{1}{4}}=\frac{1}{2}\] so \[x=\frac{7}{2}\pm \sqrt{\frac{1}{4}}=\underline{\frac{7}{2}\pm \frac{1}{2}}\] Try simplifying the underlined part

OpenStudy (ash2326):

@yomamabf did you understand?

OpenStudy (anonymous):

i already had that answer..... i needed to know why the answer was 3,4

OpenStudy (ash2326):

Yeah you just had to simplify it:)

OpenStudy (anonymous):

dude.............. fail